OpenStudy (anonymous):

a -3b = 4 a =b - 2 What is the solution to the set of equations in the form (a,b)? A. (-2,-2) B. (-3,-1) C. (-9,-7) D. (-5,-3)

OpenStudy (anonymous):

@pooja195

OpenStudy (wolf1728):

1) a -3b = 4 2) a =b - 2 Multiply equation 1) by -1 and add -b to both sides of equation 2) 1) -a +3b =- 4 2) a -b = - 2 ADD both equations 2b = -6 Can you take it from there?

OpenStudy (chrisdbest):

So A is not the answer

OpenStudy (anonymous):

i really dont know how to do it

OpenStudy (wolf1728):

For one thing what does B = ?

OpenStudy (anonymous):

3 ?

OpenStudy (chrisdbest):

D for DEEZ NUTZ!!!! GODDEIMM!!!!

OpenStudy (wolf1728):

No because 2 * 3 =6 You need a value of B so that 2 * B = -6

OpenStudy (chrisdbest):

Just Plug A In For X Values, And B In For The Y Values

OpenStudy (wolf1728):

chrisdbest what are you talking about?

OpenStudy (anonymous):

im so confused

OpenStudy (wolf1728):

itsmee you almost have the answer 2 * B = -6 what is B?

OpenStudy (wolf1728):

itsmee - how mush algebra do you know? 2 * B = -6 If you can't solve the above equation, I don't think you are ready to be solving for 2 unknowns.

OpenStudy (chrisdbest):

@wolf1728 I'm talking about the right answer without being too blunt about it.

OpenStudy (wolf1728):

I am not trying to give you a rough time. Perhaps if your instructor knew your math experience you could get into a different class.

OpenStudy (wolf1728):

Okay Chris. Sorry. I thought we weren't just supposed to "dish out" answers. It helps if someone LEARNS a little. I don't think itsmee should even be taking a class like this.

OpenStudy (wolf1728):

Okay, I'll leave and "chrisdebest" can just dish out the answers. See ya folks.